### v1

# Exponential quantum advantages in learning quantum observables from classical data

Abstract
Quantum computers are believed to bring computational advantages in simulating quantum many body systems. However, recent works have shown that classical machine learning algorithms are able to predict numerous properties of quantum systems with classical data. Despite various examples of learning tasks with provable quantum advantages being proposed, they all involve cryptographic functions and do not represent any physical scenarios encountered in laboratory settings. In this paper we prove quantum advantages for the physically relevant task of learning quantum observables from classical (measured out) data. We consider two types of observables: first we prove a learning advantage for linear combinations of Pauli strings, then we extend the result for the broader case of unitarily parametrized observables. For each type of observable we delineate the boundaries that separate physically relevant tasks which classical computers can solve using data from quantum measurements, from those where a quantum computer is still necessary for data analysis. Our results shed light on the utility of quantum computers for machine learning problems in the domain of quantum many body physics, thereby suggesting new directions where quantum learning improvements may emerge.
1 Introduction
The very first proposed application of quantum computers can be traced back to Feynman’s idea of simulating quantum physics on a quantum device. Together with factoring, simulation of quantum many body systems stand as the clearest example of dramatic advantages of quantum computers [Llo96]. Machine learning is another much newer area where quantum computers are believed to possibly bring advantages in certain learning problems, and in fact there are provable speed ups achieved by a quantum algorithm [LAT21, SG04, SSHE21, JGM+21] for specific machine learning problems. Relating back to the original Feynman’s idea of simulating quantum physics, machine learning problems in quantum many body physics seem a natural scenario where learning advantages could arise. However, perhaps surprisingly, it was recently shown that access to data exemplifying what the hard-to-compute function does can drastically change the hardness of the computational task questioning the role of quantum computation in a machine learning scenarios [HBM+21, HKT+22b]. If every quantum computation could be replicated classically provided access to data, such results would confine the practical application of quantum computers solely to the data acquisition stage. This is however not the case, as was shown already in the examples considered in [SG04, LAT21]. In such cases the unknown function was cryptographic in nature, and not related to genuine quantum simulation problems. Nonetheless in [GD23] it was elucidated that learning problems with provable speed-ups can be derived from any BQP-complete functions, facilitating connections to more physical scenarios.
The work in [GD23] demonstrated the existence of learning problems with provable advantages based on such reasonings, however it does not provide any natural learning setting of the kind encountered in physics. Another significant open question, raised by the recent progress on classical algorithms with provable guarantees, pertains to establishing clear boundaries between classical and quantum algorithms when dealing with data generated by quantum processes. In this work we provide substantial strides in both directions. We specifically consider learning problems where an experimentalist wants to predict expectation values of an unknown observable from measurements on input quantum states which can either be ground states of local Hamiltonians or time evolved states. Our result is a proof of learning advantages for different types of quantum observables. Firstly, we prove an exponential quantum advantage in learning observables constructed as an unknown linear combination of Pauli strings. Then we extend our quantum speed-up result to the far more general case of unitarily-parametrized observables, linking our results to the extensive body of literature on learning unknown unitaries from query access. In connection to this, we examine the Hamiltonian learning problem where the identification of the target function is demonstrated to be classically easy. We summarize our main results in Table 1.
Class of observables Learning Problem Existence of a Learning Advantage
Linear combination of Pauli strings Time Evolution Problem Yes
Ground State Problem Yes
Flipped Concepts No - Classically easy
Unitarily-parametrized Observables Learning the observable Yes
Hamiltonian Learning No - Classically easy
Identifying the concept Unknown - Classical hardness unknown
Table 1: Main results of this paper
2 Related work
Our main results are built upon the theoretical assumption that there exists a BQP language and an input distribution such that the language cannot be decided using $\mathrm{HeurP/poly}$, for more on this see [GD23]. Building on this, the authors also proved that many learning speed-ups could arise by looking at BQP complete problems. However, in their separation result, the concept class is limited to be of polynomial size, and the quantum algorithm crucially leverages this to find the underlying concept in polynomial time by brute force. In this work we take a step further from their result by considering a continuous concept class and using a quantum algorithm which learns the unknown concept from data through a natural machine learning method.
In recent works by [HKT+22b, LHT+24], significant progresses have been made with classical machine learning algorithms with provable guarantees for learning ground state properties from data, restricting the role of a quantum computer only to the data acquisition phase. In this study, we establish rigorous limitations on the capabilities of classical algorithms when learning quantum observables. A key distinction from their approach is that, in considering ground state problems, we investigate Hamiltonians with a polynomially decaying spectral gap, whereas the algorithms in [HKT+22b, LHT+24] have provable guarantees only for gapped Hamiltonians
In [YIM23] the authors proposed a family of supervised learning tasks which present a learning speed-up with an exponentially sized concept class. Our construction, which was developed independently from [YIM23], stems from a physically-motivated context, resulting in the utilization of different techniques for the quantum learning algorithm given the necessity to handle errors in the training samples. The result of Theorem 3 further differentiates our work, as it deals with unitarily parametrized observables for which the corresponding learning algorithm cannot be straightforwardly simplified to solving a linear system of equations, as it is in the case of [YIM23].
3 The learning problems
We now take in consideration the first scenario where experimentalists can time evolve arbitrary input states under a fixed Hamiltonian for constant time, however they do not have full control over the input state preparation. We model this by allowing the input states to be sampled randomly from an unknown distribution of quantum states, subsequently they undergo a time evolution before being measured by an unknown observable. The experimentalists collect the corresponding measurement outcomes and the goal is to predict expectation values of the unknown observable on new input states. This simple scenario is later significantly generalized to cases where the inputs are known Hamiltonians, and the problem is not time evolution but rather learning measurements on their ground states. In both cases we rigorously model the colloquially described learning tasks using concept classes within the PAC learning framework
Let $H$ be a Hamiltonian and fix a constant time $\tau$, we model the time evolution learning problem by the following concept class
$\mathcal{F}_{\mathrm{evolved}}^{H,O}$ =$=\{f^{\bm{{\bm{\alpha}}}}(\bm{x})\in\mathbb{R}\;\;|\;\mathbf{{\bm{\alpha}}}\in[-1,1]^{m}\}$ (1)
with $\text{with }\;f^{\bm{\alpha}}(\bm{x}):\;\;$ $\;\bm{x}\in\{0,1\}^{n}\rightarrow f^{\bm{\alpha}}(\bm{x})=\text{Tr}[\rho_{H}(\bm{x})O(\mathbf{{\bm{\alpha}}})]\ \;\;\&\;\;O({\bm{\alpha}})=\sum_{i=1}^{m}{\bm{\alpha}}_{i}P_{i}.$ (2)
In the above, $\bm{x}$ specifies the initial state |$\ket{\bm{x}}$, $\rho_{H}(\bm{x})$ the constant time evolved state $\rho_{H}(\bm{x})=U\ket{\bm{x}}\bra{\bm{x}}U^{\dagger}$ with $U=e^{iH\tau}$ and each $P_{i}$ is a $k$-local Pauli string where $m$ scales polynomially with $n$. Notice that although we consider binary inputs $\bm{x}\in\{0,1\}^{n}$, the output of each concept is a real. The goal of the ML learning algorithm is to learn a model $h(\bm{x})$ which approximates the unknown concept $f^{\bm{\alpha}}(\bm{x})=\text{Tr}[\rho_{H}(\bm{x})O(\mathbf{{\bm{\alpha}}})]$, using as training samples data of the form $\mathcal{T}^{\mathbf{{\bm{\alpha}}}}_{\epsilon_{3}}=\{(\bm{x}_{\ell},y_{\ell})\}_{\ell=1}^{N}$ where $y_{\ell}\approx_{\epsilon_{3}}f^{\bm{\alpha}}(\bm{x}_{\ell})$ is an additive error $\epsilon_{3}$-approximation of the true expectation value $\text{Tr}[\rho_{H}(\bm{x}_{\ell})O({\bm{\alpha}})]$. Considering datasets with approximated values makes the learning problem closer to real world scenarios. In an idealized setting, the dataset would consist of pairs ${(\bm{x}_{\ell},y_{\ell})}$, where $y_{\ell}$ represents the exact expectation value. However, in real experiments, measurements are conducted by counting frequencies from a finite number of state copies, resulting in only approximate estimations of expectation values due to sampling errors. Formally, we assume a maximum (sampling) error $\epsilon_{3}$ on the training labels $y_{\ell}$ of our dataset, i.e. $\epsilon_{3}=\max_{\ell}|\text{Tr}[\rho_{H}(\bm{x}_{\ell})O({\bm{\alpha}})]-y_{\ell}|$. It is now possible to formally state our learning condition. Assuming the $\bm{x}_{\ell}$’s in the training data come from an unknown distribution $\mathcal{D}$, we require that the ML model learns the concept class $\mathcal{F}^{H,O}_{\mathrm{evolved}}$ in the following sense.
Definition 1 (Efficient learning condition).
A concept class $\mathcal{F}$ is efficiently learnable with respect to the distribution $\mathcal{D}$ if for every $\epsilon,\epsilon_{3}>0$ and $10<\delta<$ 1, there exists a poly$(1/\epsilon,1/\delta,1/\epsilon_{3},n)$-time algorithm $A$ such that for any $f^{\bm{\alpha}}$ in $\mathcal{F}$ and for any training set $\mathcal{T}^{\bm{\alpha}}_{\epsilon_{3}}$ of poly$(1/\epsilon,1/\delta,1/\epsilon_{3},n)$ size, $h(.)=A(\mathcal{T}_{\epsilon_{3}}^{\bm{\alpha}},.)$ satisfies with probability $1-\delta$:
$\mathbb{E}_{\bm{x}\sim\mathcal{D}}\big{[}\;|f^{\bm{\alpha}}(\bm{x})-h(\bm{x})|^{2}\;\big{]}\leq\epsilon$ (3)
Notice that the learning algorithm possesses no prior knowledge regarding the observable $O({\bm{\alpha}})$, except for the fact that it is a $k$-local operator. Any additional information regarding $O({\bm{\alpha}})$ can exclusively be derived from the training samples within $\mathcal{T}^{\bm{\alpha}}_{\epsilon_{3}}$. In particular the vector ${\bm{\alpha}}$, which defines the specific concept in the concept class, is unknown to the learning algorithm. While the classical hardness of the learning problem relies on considering a specific ${\bm{\alpha}}$ for which the evaluation of $f^{\bm{\alpha}}(\bm{x})$ is known to be hard, the quantum algorithm will use a LASSO regression to infer a parameter $w\in[-1,1]^{m}$ close to the target ${\bm{\alpha}}$ so that condition (3) is satisfied.
We emphasize here that our learning definition demands that the trained classical model can label new points. This stands in contrast to other settings, e.g. Hamiltonian learning problems, where the task would be identifying the vector ${\bm{\alpha}}$. In the prior case, a learning advantage is more easily established as explained in [GD23] where the hardness of evaluating versus identifying a concept was discussed. Later, we will explore this difference further and present an example of an identification problem closely related to the concept class of Eq. (1), which indeed can be solved by a classical algorithm.
We can now state the first result of this paper, namely the existence of a concept class for the time evolution learning problem learnable by a quantum algorithm but for which no classical algorithm can meet the learning condition of Eq. (3).
Theorem 1 (Learning advantage for the time evolution problem).
Under the assumption that there exists a distribution $\mathcal{D}^{\prime}$ such that $\mathrm{BQP}\not\subset\mathrm{HeurP}^{\mathcal{D}^{\prime}}/\mathrm{poly}$ [^1] We define the class $\mathrm{HeurP}^{\mathcal{D}}/\mathrm{poly}$ as the class of languages $\mathcal{L}$ that can be decided in $\mathrm{HeurP}/\mathrm{poly}$ under the distribution $\mathcal{D}$, i.e. $\mathrm{HeurP}^{\mathcal{D}}/\mathrm{poly}=\{\mathcal{L}\;\;|\;\;(\mathcal{L},\mathcal{D})\in\mathrm{HeurP}/\mathrm{poly}\}$ , then for any $\mathrm{BQP}$-complete language there exists an associated distribution $\mathcal{D}$ over the input values $\bm{x}\in\{0,1\}^{n}$, a Hamiltonian $H_{\mathrm{hard}}$ and a set of observables $\{O({\bm{\alpha}})\}_{\bm{\alpha}}$ such that no classical algorithm can solve the time evolution learning problem, formalized by the concept class $\mathcal{F}_{\mathrm{evolved}}^{H_{\mathrm{hard}},O}$, in the sense of Def. 1 given data $\mathcal{T}^{\mathbf{{\bm{\alpha}}}}_{\epsilon_{3}}=\{(\bm{x}_{\ell},y_{\ell}\approx_{\epsilon_{3}}f^{\bm{\alpha}}(\bm{x}))\}_{\ell=1}^{N}$ sampled from $\mathcal{D}$. However, there exists a quantum algorithm which learns $\mathcal{F}_{\mathrm{evolved}}^{H_{\mathrm{hard}},O}$ under any distribution $\mathcal{D}$.
In the remainder of this Section we prove Theorem 1 by explicitly constructing an example of a provable classically hard time evolution problem and providing a quantum algorithm with learning guarantees. We sketch the proof ideas in the following paragraphs of the main text, while the full proofs can be found in the Appendix.
Classical hardness The hardness condition relies on the assumption $\mathrm{BQP}\not\subset\mathrm{HeurP}^{\mathcal{D}}/\mathrm{poly}$ for some distribution $\mathcal{D}$. Notice that it is a widely believed assumption as, for example, the discrete logarithm problem is believed to not be in $\mathrm{HeurBPP}/\mathrm{poly}$ if the input values are sampled from the uniform distribution [BM19]. In general, by Lemma 3 in [GD23] if there exists a single $\mathcal{L}\in\text{BQP}$ that is not in HeurP/poly under some distribution, then for every BQP-complete problem there exists a distribution under which the problem is not in HeurP/poly.
Lemma 1 (Classical hardness of the time evolution learning problem).
Under the assumption that there exists a distribution $\mathcal{D}^{\prime}$ such that $\mathrm{BQP}\not\subset\mathrm{HeurP}^{\mathcal{D}^{\prime}}/\mathrm{poly}$, then for any BQP-complete language there exists an associated input distribution $\mathcal{D}$ and a Hamiltonian $H_{\mathrm{hard}}$ such that no randomized polynomial-time classical algorithm $A_{c}$ is such that $h_{cl}(.)=A_{c}(T^{\bm{\alpha}},.)$ satisfies the learning condition of Def. 1 for the concept class $\mathcal{F}^{H_{\mathrm{hard}},O}_{\mathrm{evolved}}$ with $x$ sampled from $\mathcal{D}$.
We now show an appropriate choice of the Hamiltonian $H_{\mathrm{hard}}$ such that the time evolution problem lies outside $\mathrm{HeurP}/\mathrm{poly}$ for a specific input distribution $\mathcal{D}$. Let $\mathcal{L}$ be a $\mathrm{BQP}$-complete language and $\mathcal{D}$ a distribution over input values such that $(\mathcal{L},\mathcal{D})\not\subset\mathrm{HeurP}/\mathrm{poly}$. Since $\mathcal{L}\in\mathrm{BQP}$ there exists a quantum circuit $U_{BQP}$ which decides input bitstrings $x\in\{0,1\}^{n}$ correctly on average with respect to $\mathcal{D}$. As shown more rigorously in the Appendix, measuring the $Z$ operator on the first qubit of the state $U_{BQP}\ket{\bm{x}}$ will output a positive or negative value depending if $x\in\mathcal{L}$ or not. Therefore, considering the observable $O^{\prime}=Z\otimes I\otimes...\otimes I$ the quantum model $f^{O^{\prime}}(\bm{x})=\text{Tr}[O^{\prime}\rho_{U_{BQP}}(\bm{x})]$, with $\rho_{U_{BQP}}(\bm{x})=U_{BQP}\ket{\bm{x}}\bra{\bm{x}}U^{\dagger}_{BQP}$, correctly decides every input $\bm{x}$. We then show in the Appendix that if a classical algorithm can learn the function $g(\bm{x})=\text{Tr}[O^{\prime}\rho_{U_{BQP}}(\bm{x})]$ up to the learning condition in Def. 1, then it would be able to heuristically decide the language $\mathcal{L}$ with respect to the distribution $\mathcal{D}$. Finally, using Feynman’s idea [Fey85, Nag10] it is possible to construct a local Hamiltonian which time-evolves the initial state |$\ket{\bm{x}}$ into $U_{BQP}\ket{\bm{x}}$ in constant time [Chi04]. Considering as $H_{\mathrm{hard}}$ such constructed Hamiltonian concludes our proof. In fact, since there exists at least one concept (namely the one labeled by ${\bm{\alpha}}^{\prime}$ such that $O({\bm{\alpha}}^{\prime})=O^{\prime}$) which is not learnable by a classical algorithm, then we can conclude that for such kind of Hamiltonians the time evolution problem is not classically learnable.
Quantum learnability To establish the quantum learnability of the classically hard concept class constructed above for the time evolution problem, we present a quantum algorithm directly. This algorithm accurately predicts any concept in the associated learning problem with high probability when provided with a sufficient number of training data samples. Crucially, we prove that the algorithm meets the learning condition of Definition 1 using only polynomial training samples and running in polynomial time. The central idea involves leveraging the capability of the quantum algorithm to efficiently prepare the time-evolved states $\rho_{H(\bm{x})}$, for any input local Hamiltonians $H$ and, in particular, for the hard instances of $H_{\mathrm{hard}}$ considered in our hardness result of Lemma 1.
Lemma 2 (Quantum learnability of the time evolution learning problem).
There exist a quantum algorithm $A_{q}$ such that for any concept $f^{\bm{\alpha}}\in\mathcal{F}^{H_{\mathrm{hard}},O}_{\mathrm{evolved}}$ considered in Lemma 1, $A_{q}$ satisfies the following. Given $n,\epsilon,\delta\geq 0$, and any training dataset $T^{{\bm{\alpha}}}_{\epsilon_{3}}$ of size
$N=\mathcal{O}\left(\frac{\log(n/\delta)n}{\epsilon}\right)$ (4)
with probability $1-\delta$ the quantum algorithm $A_{q}$ outputs a model $h^{*}(.)=A(T^{{\bm{\alpha}}}_{\epsilon_{3}},.)$ which satisfies the learning condition of Def. 1:
$\mathbb{E}_{\mathbf{x}\sim\mathcal{D}(\mathbf{x})}\big{[}\;|f^{{\bm{\alpha}}}(\mathbf{x})-h^{*}(\bm{x})|^{2}\;\big{]}\leq\epsilon$ (5)
where $\mathcal{D}$ is the distribution the training data is sampled from.
The idea for the quantum algorithm is the following. For every point $\bm{x}_{\ell}$ we construct a vector $\phi(\bm{x}_{\ell})\in[-1,1]^{m}$ of expectation values of the single Pauli strings present in $O$ on the time evolved quantum states. The model $h(\bm{x})=w\cdot\phi(\bm{x})$ is then trained on the data samples with a LASSO regression to find a $w^{*}$ so that the trained model is in agreement with the training samples, i.e. $h^{*}(\bm{x}_{\ell})=w^{*}\cdot\phi(\bm{x}_{\ell})\approx y_{\ell}$ for any $(\bm{x}_{\ell},y_{\ell})\sim\mathcal{T}^{\bm{\alpha}}_{\epsilon_{3}}$. As the $\ell_{1}$-norm of the optimal $w_{opt}={\bm{\alpha}}$ scales polynomially in $n$, imposing the constraint $||w||_{\ell_{1}}\leq B$ with $B=\mathcal{O}(n)$ in the LASSO regression will allow to obtain an error $\epsilon$ in the generalization performance using a training set of at most polynomial size. This is because the generalization error for the LASSO regression is bound linearly by $B$ [MRT18]. We write here the description of the quantum algorithm, while we leave the precise analysis of its sample and time complexity in the Appendix.
Algorithm 1 Quantum Algorithm
  1. 1.
    For every training point in $T^{\mathbf{{\bm{\alpha}}}}_{\epsilon_{3}}=\{(\bm{x}_{\ell},y_{\ell})\}_{i=1}^{N}$ the quantum algorithm prepares poly($n$) copies of the state $\rho_{H}(\bm{x}_{\ell})$ and computes the estimates of the expectation values $\langle P_{j}\rangle_{\ell}=\text{Tr}[\rho_{H}(\bm{x}_{\ell})P_{j}]\;\;\forall j=1,...,m$ up to a certain precision $\epsilon_{1}$. Note that $m$ scales at most polynomially in $n$ as $\{P_{j}\}_{j=1}^{m}$ are local observables.
  2. 2.
    Define the model $h(\bm{x})=w\cdot\phi(\bm{x})$, where $\phi(\bm{x})$ is the vector of the Pauli string expectation values $\phi(\bm{x})=\left[\text{Tr}[\rho_{H}(\bm{x})P_{1}],...,\text{Tr}[\rho_{H}(\bm{x})P_{m}]\right]$ computed at Step 1. Then given as hyperparameter a $B\geq 0$ the LASSO ML model finds an optimal $w^{*}$ from the following optimization process:
    $\min_{\begin{subarray}{c}w\in\mathbb{R}^{m}\\ ||w||_{1}\leq B\end{subarray}}\;\;\frac{1}{N}\sum_{l=1}^{N}|w\cdot\phi(\bm{x}_{l})-y_{l}|^{2}$ (6)
    with $\{(\bm{x}_{l},y_{l}=\text{Tr}[\rho_{H}(\bm{x}_{l})O({\bm{\alpha}})])\}_{l=1}^{N}$ being the training data.
    Importantly, to meet the learning condition the optimization does not need to be solved exactly, i.e. $w^{*}={\bm{\alpha}}$. As we will make it clear in the Appendix, it is sufficient to obtain a $w^{*}$ whose training error is $\epsilon_{2}$ larger than the optimal one.
3.1 Generalizations: quantum advantages for fixed inputs and for ground state problems
We showed that the concept class defined in Eq. (1) leads to a learning advantage. It is important to note that, unlike many other common learning scenario settings, the physical problem modeled by such concept class implies that the experimentalists cannot select the initial input states. However, they do possess precise control over the Hamiltonian governing the evolution of these input states. It is natural to ask if we can generalize the result and provide other cases of exponential learning advantages where the experimentalist has instead only little control over the Hamiltonian employed.
Fixed inputs, unknown Hamiltonians A first example that naturally arises from the case discussed above is when the experimentalists prepare a fixed initial state and time evolves it under a Hamiltonian within a fixed family of Hamiltonians but over which they do not have control. Specifically the Hamiltonians in such a family will be labeled by some input bitstring $\bm{x}$, which for example could parameterize the strength of the coupling interactions. Each initial state will then be evolved by a different Hamiltonian in the family accordingly to the input $\bm{x}$, which comes from an unknown underlying distribution. It is easy to see that the mathematical description of such a learning problem is again defined by the concept class in Eq.(1). The only change is in the definition of $\rho_{H}(\bm{x})$, now being $\rho_{H}(\bm{x})=U(\bm{x})\ket{0}\bra{0}U(\bm{x})^{\dagger}$ with $H(\bm{x})$ the Hamiltonian associated to the quantum circuit $U(\bm{x})$ by the Feynmann construction [Fey85]. A learning advantage exists for such concept class as well, as the general definition of a language $\mathcal{L}$ in BQP implies the existence of a family of circuits $\{U(\bm{x})\}_{x}$ which correctly decides every $\bm{x}\in\mathcal{L}$.
Ground state problem As our next example we consider the task of predicting ground state properties of local Hamiltonians. Here, the states to be measured are no longer time evolved quantum states but rather ground states of local Hamiltonians. Specifically, the experimentalist is capable of preparing ground states of input $k$ local Hamiltonians, which once again belong to a family of Hamiltonians. However, they lack control over the coupling parameters that, instead, are random values drawn from an underlying distribution. This situation occurs, for example, in the context of the random Ising [Nat98] or random Heisenberg model [OS01]. The mathematical formalization of such a learning problem is the following. Consider a family of parametrized local Hamiltonians =$\mathcal{H}=\{H(\bm{x})\;\;|\;\;\bm{x}\in\{0,1\}^{n}\}$. Let us define the concept class for the ground state learning problem similarly to the time evolution case of Eq. (1), where now the unknown observable $O({\bm{\alpha}})$ is measured on the states $\rho_{H}(\bm{x})$, which correspond to the ground states of the Hamiltonians $H(\bm{x})\in\mathcal{H}$, then we define:
$\mathcal{F}^{\mathcal{H},O}_{\mathrm{g.s.}}$ =$=\{f^{\bm{\alpha}}(\bm{x})\in\mathbb{R}\;\;|\;\mathbf{{\bm{\alpha}}}\in[-1,1]^{m}\}$ (7)
with $\text{with }f^{\bm{\alpha}}(\bm{x}):\;\;$ $\;\bm{x}\in\{0,1\}^{n}\rightarrow f^{\bm{\alpha}}(\bm{x})=\text{Tr}[\rho_{H}(\bm{x})O(\mathbf{{\bm{\alpha}}})]\ \;\;\&\;\;O({\bm{\alpha}})=\sum_{i=1}^{m}{\bm{\alpha}}_{i}P_{i}.$ (8)
Considering the training data $\mathcal{T}^{\mathbf{{\bm{\alpha}}}}_{\epsilon_{3}}=\{(\bm{x}_{\ell},y_{\ell}\approx_{\epsilon_{3}}f^{\bm{\alpha}}(\bm{x}_{\ell}))\}_{\ell=1}^{N}$, the learning condition remains the same as in Definition 1. From the hardness result of Lemma 1, we obtain the following Theorem
Theorem 2 (Learning advantage for the ground state learning problem).
Under the assumption that there exists a distribution $\mathcal{D}^{\prime}$ such that $\mathrm{BQP}\not\subset\mathrm{HeurP}^{\mathcal{D}^{\prime}}/\mathrm{poly}$, then for any BQP-complete language there exists an associated distribution $\mathcal{D}$ over the input values $x\in\{0,1\}^{n}$ and a family of Hamiltonians $\mathcal{H}_{\mathrm{hard}}$ such that no classical algorithm can learn the ground state problem, formalized by learning the concept class $\mathcal{F}^{\mathcal{H}_{\mathrm{hard}},O}_{\mathrm{g.s.}}$ in the sense of Def.1 given data $\mathcal{T}^{\mathbf{{\bm{\alpha}}}}_{\epsilon_{3}}=\{(\bm{x}_{\ell},y_{\ell}\approx f^{\bm{\alpha}}(\bm{x}))\}_{\ell=1}^{N}$ sampled from $\mathcal{D}$. However, there exists a quantum algorithm which learns $\mathcal{F}^{\mathcal{H}_{\mathrm{hard}},O}_{\mathrm{g.s.}}$ under any distribution $\mathcal{D}$.
Proof.
The existence of a class of Hamiltonian $\mathcal{H}_{\mathrm{hard}}$ for which the ground state problem is classically hard to learn is guaranteed by the argument above regarding the hardness of time evolution case. The structure of the proof is exactly the same of the proof for Theorem 1 rigorously written in the Appendix. The only missing step for the ground state version is that now the states $\rho_{H}(\bm{x})$ we are considering in Eq. (7) are ground state of Hamiltonians $H(\bm{x})$ and not time evolved states. However, using the Kitaev’s construction [KSV02, KR03], it is possible to create for any $\bm{x}\in\{0,1\}^{n}$ and $U$ a local Hamiltonian $H(\bm{x})$ such that its ground state will have a large overlap with $U\ket{\bm{x}}$, with $U$ an arbitrary quantum circuit with a polynomial depth. This completes our proof of classical hardness as we consider the family of Hamiltonian $\mathcal{H}_{\mathrm{hard}}$ to exactly be the set of such $\{H(\bm{x})\}_{x}$ with $U$ implementing a BQP-complete computation [^2] More rigorously it means that for every $n$, we consider $U=U^{n}_{BQP}$ from the family $\{U^{n}_{BQP}\}_{n}$ of quantum circuits which correctly decide a BQP-complete language $\mathcal{L}$. Same as before, it is still the case that there is at least one concept which can not be evaluated by polynomially sized classical circuits. Finally, also the quantum algorithm with learning guarantees for the concept class $\mathcal{F}^{\mathcal{H}_{\mathrm{hard}},O}_{\mathrm{g.s.}}$ closely follows Algorithm 1. The only missing point to prove here is that the ground states $\rho_{\mathcal{H}_{\mathrm{hard}}}(\bm{x})$ are easily preparable on a quantum computer. Recall that the class of hard Hamiltonian $\mathcal{H}_{\mathrm{hard}}$ considered in the hardness result of the ground state problem are the one derived from the Kitaev’s circuit-to-Hamiltonian construction from a $\mathrm{BQP}$-complete circuit. It is well known that those Hamiltonians present a $\Omega(1/\text{poly}(n))$ gap (in contrast to the classically learnable Hamiltonians in [HKT+22b]) and it is possible to construct the ground state $\rho_{\mathcal{H}_{\mathrm{hard}}}$ from the description of the corresponding Kitaev’s Hamiltonian, known to the learner through the description of the concept class and the input $\bm{x}$. This then concludes our proof for the quantum learnability of $\mathcal{F}^{\mathcal{H}_{\mathrm{hard}},O}_{\mathrm{g.s.}}$, thus completing the entire proof of Theorem 2
Notice that if our interest is only about the classical hardness or the quantum learnability of the ground state problem we can consider concept classes associated to more general Hamiltonian families than the Kitaev’s family $\mathcal{H}_{\mathrm{hard}}$ used in our provable advantage result. Specifically, classical hardness for the time evolution problem is guaranteed for a number of physically motivated Hamiltonians, such as the Heisenberg model whose dynamics can efficiently be simulated on a quantum computer. Notice that for the ground state version of the learning problem considering generic local Hamiltonian might make the problem quantum not solvable. It is in fact known that finding ground states of even two local Hamiltonians, as for example the Heisenberg model, is QMA-hard. Concept classes for which the quantum learnability condition is met are the ones associated to any family of Hamiltonian $\mathcal{H}_{\mathrm{quantum}}$ for which a quantum algorithm can prepare the corresponding ground state given the Hamiltonian description. In that regard a learning speed-up can still be achieved if we consider providing to the algorithm guiding states for any family of local Hamiltonians, this in fact will make the problem $\mathrm{BQP}$-complete [GHGM22].
We note that the ground state learning problem defined by the class in Eq.(7) very much resembles the ML problem studied in the works [HKT+22b] and [LHT+24] where the authors showed that a classical algorithm could solve the task. As their result holds for any local Hamiltonian with constant gap, this gives a constraint on the family of local Hamiltonians $\mathcal{H}$ one would like to consider in order to prove a quantum advantage in learning.
The “flipped case” As a final remark, it is interesting to observe that if we consider the case of fixed input state, which could be a fixed initial state $\rho_{0}(\bm{x})=\ket{\bm{x}}\bra{\bm{x}}$ | in the time evolution scenario modeled by the concept class in Eq. (1) or a single ground state $\rho_{H}(\bm{x})$ of a fixed Hamiltonian $H(\bm{x})$ in the ground state problem, then the learning problem becomes trivially classically easy. Such learning scenario is formally equivalent to considering a “flip concept” of the ones defined above. Let us for example consider the task of learning an unknown quantum process from many measurements, in this case the learning problem is still modeled by the concept class $\mathcal{F}^{H,O}_{\mathrm{evolved}}$ of Eq. (1) with the distinction that now the role of $x$ and ${\bm{\alpha}}$ are switched. Namely, the concepts are defined as $f^{\bm{x}}({\bm{\alpha}})$ so that the their expressions remain the same of $f^{\bm{\alpha}}(\bm{x})$. The difference lies in the labeling, where $x$ now denotes the concept while ${\bm{\alpha}}$ represents the input vectors. As $\bm{x}$ is constant for an instance of the learning problem, specified by the concept that generates the data, the training samples are measurements of the same quantum state $\rho_{H(\bm{x})}$ with different observables corresponding to different ${\bm{\alpha}}$. Since $O({\bm{\alpha}})=\sum_{j}{\bm{\alpha}}_{j}P_{j}$, we can make use of data to solve the linear system and obtain the expectation values of each local Pauli string $\text{Tr}[\rho_{H_{\mathrm{hard}}}(\bm{x})P_{i}]$. It then becomes easy to extrapolate the value of $\text{Tr}[\rho_{H_{\mathrm{hard}}}(\bm{x})O({\bm{\alpha}})]$ for every new ${\bm{\alpha}}$.
3.2 Generalization of the quantum learning mechanism to quantum kernels
In [LAT21], considerable effort was invested to construct a task with provable learning speed-up where the quantum learning model is somewhat generic (related to natural QML models studied in literature), while still being capable of learning a classically unlearnable task. In our construction, it is possible to view the entire learning process as a quantum kernel setting, similarly to the approach in [YIM23]. However, in standard kernel approaches, especially those stemming from support vector machines, the optimization process is solved in the dual formulation where the hypothesis function is expressed as a linear combination of kernel functions. In contrast, the LASSO optimization employed here solves the optimization in the primal form with a constrain on the $\ell_{1}$ norm. This is an issue making our learner not technically a quantum kernel. While the LASSO formulation does not straightforwardly convert into a kernel method, one could attempt to address the regression problem outlined in Equation (6) using an alternative optimization approach that supports a kernel solution, such as kernel ridge regression. In this case the optimization is done for vectors with bounded $\ell_{2}$ norms, nevertheless there exist bounds on the generalization performance of such procedures as well [MRT18].
4 Generalization to observables parametrized by unitaries
We have demonstrated the existence of a quantum advantage for the learning problem of predicting $k$-local observables from time evolved states and from ground states. Critically, we considered observables of the type $O({\bm{\alpha}})=\sum_{i}{\bm{\alpha}}_{i}P_{i}$ and we exploited the linear structure of $O({\bm{\alpha}})$ to ensure quantum learnability through LASSO regression. It is however natural to ask if quantum learnability can be achieved for other types of observables, while maintaining the classical hardness. In this section we consider the far more general case where the unknown observable is parameterized through a unitary matrix, i.e. $O({\bm{\alpha}})=W({\bm{\alpha}})OW^{\dagger}({\bm{\alpha}})$ where $O$ is an hermitian matrix. Our findings in this scenario will be of two kinds. First we show as a general result that for every method which learns a unitary $W$ given query access on a known distribution of input quantum states there exists a learning problem which exhibits a classical-quantum advantage. Then we concretize the general result by presenting a constructed example of a learning problem defined by a class of unitary-parameterized observables, showcasing a provable speed-up. Before stating our findings, let us properly introduce the learning problem under consideration. Imagine the scenario where the experimentalist once more measures observables on evolved quantum states. However, in this case, they lack control over the entire evolution, with a portion of it remaining unknown. Specifically, the experimentalist will receive measurements on input-dependent states |$\ket{\psi(\bm{x})}=W({\bm{\alpha}})U(\bm{x})\ket{0}$ for an unknown fixed ${\bm{\alpha}}$. Their objective remains predicting expectation values on states |$\ket{\psi(\bm{x}^{\prime})}$ for new inputs $\bm{x}^{\prime}$. Concretely such scenario corresponds to the following concept class:
$\mathcal{M}_{U,W,O}$ =$=\{f^{\bm{\alpha}}(\bm{x})\in\mathbb{R}\;\;|\;\;{\bm{\alpha}}\in[-1,1]\}$ (9)
with $\text{with }\;\;f^{\bm{\alpha}}(\bm{x}):\;x$ $\in\mathcal{X}\subseteq\{0,1\}^{n}\rightarrow\text{Tr}[\rho_{U}(\bm{x})O({\bm{\alpha}})]\in\mathbb{R}\;\;\&\;\;O({\bm{\alpha}})=W({\bm{\alpha}})OW^{\dagger}({\bm{\alpha}}).$ (10)
where $\rho_{U}(\bm{x})=U(\bm{x})\ket{0}\bra{0}U^{\dagger}(\bm{x})$ and $O$ is a hermitian matrix. Given as training data $\mathcal{T}^{\bm{\alpha}}=\{(\bm{x}_{\ell},y_{\ell})\}_{\ell=1}^{N}$ with $\mathbb{E}[y_{\ell}]=\text{Tr}[\rho_{U}(\bm{x}_{\ell})O({\bm{\alpha}})]$, the goal of the learning algorithm is again to satisfy the learning condition of Def. 1. We are now ready to state our main general result of this Section:
Theorem 3 (Learning advantage for unitarily-parametrized observables).
Every (non-adaptive) [^3] Non-adaptive means that the algorithm probes the unitary with some input states and measures some observables which are chosen independently, meaning they do not have any choice which depends on previous outcomes. learning algorithm $\mathcal{A}_{W}$ for learning a unitary $W(\bm{\alpha})\in\{W(\bm{\alpha})\}_{\bm{\alpha}}$, where the probe states {|$\{\ket{\psi_{l}}\}_{l}$and observables $\{Q_{m}\}_{m}$ come from discrete sets $S=\{\ket{\psi_{\ell}}\}_{\ell}$ and $Q=\{Q_{m}\}_{m}$ (or they can be discretized with controllable error), induces a classical-input-classical-output learning problem with a quantum-classical learning advantage.
In other words, Theorem 3 guarantees that whenever there is an efficient method to learn a unitary from query access on a set of arbitrary input states and observables it is also possible to construct a learning problem of the kind Eq. (9) which exhibits a learning speed-up. We will more mathematically formalize the statement of Theorem 3 and provide a rigorous proof in the Appendix. The nontrivial part of our result lies in the fact that in the majority of the works present in the literature, which provide efficient algorithms to learn a unitary, the set of required probe states {|$\{\ket{\psi_{\ell}}\}_{\ell}$ is restricted to specific quantum states, often classically describable (e.g. stabilizer states). On such a set of states, our hardness results from the previous section can not directly be applied as a classical learning algorithm could simply prepare those states as the quantum algorithm would do.
The detailed proof of Theorem 3 can be found in the Appendix, here we outline the main ideas.
Proof sketch.
The main idea of the proof is to construct a concept class of the kind of Eq.(9) which can be learned by a quantum algorithm using the learning algorithm $\mathcal{A}_{W}$, while maintaining classical harndess. Consider the circuit in Figure 1. Let $U(\bm{x})$ be a unitary that, depending on the first bit $x_{1}$ of the input $\bm{x}\in\{0,1\}^{n}$, prepares the |$\ket{\psi^{n_{S}}_{U}(\bm{x})}$ on the $n_{S}$ qubit register in two different ways. If $x_{1}=0$, then |$\ket{\psi^{n_{S}}_{U}(\bm{x})}$ is exactly one of the states |$\ket{\psi_{\ell}}\in S$, labeled by the final $n_{S}$ bits $\bm{x_{S}}\in\{0,1\}^{n_{S}}$ of the input bitstring $\bm{x}$, i.e. $\bm{x_{S}}=x_{n-n_{S}}...x_{n}$. If $x_{1}=1$, then $U(\bm{x})$ prepares the state |$\ket{\psi^{n_{S}}_{U}(\bm{x})}$ as the result of a BQP-complete computation. Regarding the observable, we define a controlled unitary $V_{A}$, controlled by the first $1+n_{Q}$ qubits register, such that when $x_{1}=0$ $V_{A}$ acts on the target $n_{S}$ qubit register by rotating the $n_{S}$ qubit measurement operator $O$ into one of the $Q_{m}$ in the set $Q$ required by the learning algorithm. The description of which $Q_{m}$ the unitary $V_{A}$ implements is contained in the $n_{Q}$ bitstring $\bm{x_{Q}}\in\{0,1\}^{n_{Q}}$ with $\bm{x_{Q}}=x_{2}x_{3}...x_{n_{Q}+1}$. When $x_{1}=1$, $V_{A}$ acts as the identity matrix on the $n_{S}$ qubit register. Suppose now the input bits are sampled from the following distribution $\mathcal{D}$ on $\bm{x}\in\{0,1\}^{n}$:
  • The first bit $x_{1}$ of $\bm{x}$ is randomly selected with equal probability between 0 and 1.
  • If $x_{1}=0$ then the other $n-1$ bit $x_{2}x_{3}...x_{n}$ are sampled from the required distribution [^4] Note that the distribution of probes states and measurements required by $\mathcal{A}_{W}$ to learn the unitary $W(\bm{\alpha})$ could be very complicated involving perhaps a joint distribution on probe states and measurements. Nevertheless every target distribution can be obtained by post-processing of samples from the uniform distribution, which can be done coherently by taking uniformly random input bitstrings. by $\mathcal{A}_{W}$ to learn the unitary $W(\bm{\alpha})$.
  • If $x_{1}=1$ then the $n_{S}$ bits in $\bm{x_{S}}$ are sampled according to a hard distribution which makes the BQP computation of $U(\bm{x})$ classically hard on average. As discussed before, such distribution always exists for BQP-complete problems assuming the existence of an input distribution $\mathrm{BQP}\not\subset\mathrm{HeurP}^{\mathcal{D}^{\prime}}/\mathrm{poly}$.
Taking $W(\bm{0})=I^{\otimes n_{S}}$ for $\bm{\alpha}=\bm{0}$ and $O=I\otimes I^{\otimes n_{Q}}\otimes Z\otimes I^{n_{S}-1}$, it is clear that the concept $f^{0}=\text{Tr}[\rho_{U}(\mathbf{x})V_{A}OV^{\dagger}_{A}]$ cannot be learned by any classical algorithm. In particular, by the result of our Lemma 1 any classical algorithm can not meet the learning condition of Def.(1) on more than half of the inputs $\bm{x}$ coming from $\mathcal{D}$, namely it is guaranteed to fail for the inputs for which $x_{1}=0$. To prove quantum learnability we notice that for every $\bm{\alpha}$ the dataset $\mathcal{T}^{{\bm{\alpha}}}$ associated to the concept class $\mathcal{M}_{U,W,O}$ contains exactly the pairs of state and measurement outcomes needed by the learning algorithm $\mathcal{A}_{W}$ in order to learn the unitary $W(\bm{\alpha})$. Namely, half of the training samples, characterized by $x_{1}=0$ in their input $\bm{x}$, allows the algorithm to recover the unknown $W(\bm{\alpha})$. Therefore the quantum algorithm is able to learn the unknown observable $O({\bm{\alpha}})$ and evaluate it on each quantum state associated to every input $\bm{x}\in\{0,1\}^{n}$. ∎
As a corollary of Theorem 3, we construct a concrete example of learning advantage for observables parametrized by shallow unitaries by exploiting the recent results on learning shallow unitaries in [HLB+24].
Corollary 1 (Learning advantage for shallow unitaries).
There exists a family of parametrized unitaries $\{U(\bm{x})\}_{x}$ and parametrized shallow circuits $\{W({\bm{\alpha}})\}_{\bm{\alpha}}$, a measurement $O$ and a distribution $\mathcal{D}$ over $x\in\{0,1\}^{n}$ such that the concept class $\mathcal{M}_{U,W,O}$ is not classically learnable with respect to the input distribution $\mathcal{D}$. However, there exists a quantum algorithm which learns $\mathcal{M}_{U,W,O}$ on the input distribution $\mathcal{D}$.
The full proof of Corollary 1 can be found in the Appendix.
{quantikz}
[…]
4.1 Relationship to Hamiltonian learning
In order to further elucidate the lines between settings with and without classical/quantum learning advantages, we can establish a connection between the task of learning Hamiltonians from time-evolved quantum states and learning unitarily-parametrized observables. The setting of the Hamiltonian learning problem is the following. Given an unknown local Hamiltonian in the form of $H(\bm{\lambda})=\sum_{i}\lambda_{i}P_{i}$, the objective of the Hamiltonian learning procedure is to recover $\bm{\lambda}$. In the version of the problem we consider, we are given a black box which implements the time evolution under the unknown Hamiltonian $H(\bm{\lambda})$ on any arbitrary quantum state. The black box action on arbitrary inputs $\rho$ is given by
$\rho\rightarrow U\rho U^{\dagger}$ (11)
with $U=e^{itH(\bm{\lambda})}$, where the evolution time $t$ is known to us. In [HKT22a] the authors provide a classical algorithm which learns the unknown $\bm{\lambda}$ from the expectation values of the Pauli string $P_{i}$ on polynomially many copies of the evolved state $U\rho U^{\dagger}$, for particular choices of initial states $\rho$’s. We can rephrase the Hamiltonian learning task in terms of a learning problem with many similarities to the ones we considered before. Concretely we define the corresponding concept class
$\mathcal{F}_{\bm{\lambda}}$ =$=\{f^{\bm{\lambda}}(\bm{x})\in\mathbb{R}\;\;|\;\;\bm{\lambda}\in[-1,1]\}$ (12)
with $\text{with }\;f^{\bm{\lambda}}(\bm{x}):$ $\;x\in\mathcal{X}\subseteq\{0,1\}^{n}\rightarrow\text{Tr}[\rho(\bm{x})O(\bm{\lambda})]\;\;\&\;\;O(\bm{\lambda})=\sum_{i}^{m}U(\bm{\lambda})^{\dagger}P_{i}U(\bm{\lambda}).$ (13)
where $x$ describes the input state $\rho(\bm{x})$, and $U=e^{itH(\bm{\lambda})}$. Considering $\mathcal{T}^{\bm{\lambda}}={(\bm{x}_{\ell},\text{Tr}[\rho(\bm{x}_{\ell})O(\bm{\lambda})])}_{\ell}$ as training data, the connection to the Hamiltonian learning task of [HKT22a] becomes evident if we demand that the learning algorithm needs only to identify, rather than evaluate, the correct concept that generated the data. It is worth noting that the concept class $\mathcal{F}_{\bm{\lambda}}$ closely resembles the concept class for the time evolution problem described earlier, for which a learning advantage was demonstrated. The main difference is that in $\mathcal{F}_{\bm{\lambda}}$ the unknown observable is no longer a linear combination of Pauli strings, instead each Pauli string in $O(\bm{\lambda})$ is parameterized by a unitary. As we are requiring to only identify the correct concept and considering that the authors in [HKT22a] provide an algorithm capable of solving the Hamiltonian Learning problem in polynomial time (for sufficiently small $t$), one might suppose that the classical difficulty associated with the learning advantage presented in this work arises solely from the challenge of evaluating the correct concept rather than from the task of identifying it. This would answer another question left open in [GD23], at least in the case of concepts labeled by observables with an unknown unitary parameterization. However it must be noted that for the Hamiltonian learning algorithm in [HKT22a] to work, the input states $\rho(\bm{x})$ in Eq.(12) are of a very particular and simple form and do not come from a distribution $\mathcal{D}$ of $\mathrm{BQP}$ complete quantum states $\rho_{H_{\mathrm{hard}}}(\bm{x})$ as we assumed in our classical hardness result. For general input distributions, the hardness of identification thus remains elusive. In this regard it is important to note that the task of identification can take various forms. For instance, here we are focusing on identifying the concept within the concept class that generated the data, and in this case, it is generally unclear when this setting allows a classical-efficient solution. In a more general scenario one could ask if the learning algorithm is able to identify a model from a hypothesis class that differs from the concept class but that it is still capable of achieving the learning condition. Colloquially here the question becomes: can traditional machine learning determine which quantum circuit could accurately label the data, even though classical computers lack the capability to evaluate such quantum circuits. In [GD23] it was proven that in this broader context provable speed-ups in identification are no longer possible. This is because the classical algorithm can successfully identify a sophisticated quantum circuit that carries out the entire quantum machine learning process as its subroutine, thereby delegating the learning aspect to the labeling function itself, prior to evaluating a data point. Thus the task of identification is indeed only interesting when the hypothesis class is somehow restricted, which is often the case in practically relevant learning scenarios.
Acknowledgments
This publication is part of the project Divide & Quantum (with project number 1389.20.241) of the research programme NWA-ORC which is (partly) financed by the Dutch Research Council (NWO). This work was also supported by the Dutch National Growth Fund (NGF), as part of the Quantum Delta NL programme. This work was also supported by the European Union’s Horizon Europe program through the ERC CoG BeMAIQuantum (Grant No. 101124342)
Appendix A Proof of Theorem 1
As anticipated in the main text, the proof of the main theorem of our work consists of two parts. First, we show that there exists a family of hamiltonians $H_{\mathrm{hard}}\in\mathcal{H}_{\mathrm{hard}}$ and observables $O$ such that the concept class $\mathcal{F}^{H_{\mathrm{hard}},O}_{\mathrm{evolved}}$ is not classically learnable, unless $\mathrm{BQP}\subseteq\mathrm{HeurP}^{\mathcal{D}}/\mathrm{poly}$ for every distribution $\mathcal{D}$. Second, we rigorously prove that Algorithm 1 learns $\mathcal{F}^{H_{\mathrm{hard}},O}_{\mathrm{evolved}}$ in polynomial time. We formally derive the proofs for the concept class $\mathcal{F}^{H_{\mathrm{hard}},O}_{\mathrm{evolved}}$ of measurements on time-evolved states. However, as we discussed in the main text, the same proofs can be easily extended for the class $\mathcal{F}^{\mathcal{H},O}_{\mathrm{g.s.}}$ of measurements on ground state of local Hamiltonians through the Kitaev’s circuit-to-hamiltonian construction. Before stating the main proof, let prove an useful lemma which guarantees the BQP-completeness of constant time Hamiltonian evolution following the idea in present in [Chi04].
Lemma 3 (Constant time evolution is BQP-complete).
For any $k-$gate quantum circuit $U=U_{k}...U_{2}U_{1}$ which acts on $n$ qubits there exists a local Hamiltonian $H$ such that for any $n$ qubit inital state |$\ket{\psi}$ :
[…]
for $t=\pi$.
Proof.
Consider the Feymann clock Hamiltonian [Fey85, Nag10]:
$H:=\sum_{j=1}^{k}H_{j}$ (15)
where:
$H_{j}=U_{j}\otimes\ket{j}\bra{j-1}+U^{\dagger}_{j}\otimes\ket{j-1}\bra{j}$ | (16)
[…]
Notice first that in the subspace spanned by the vectors {|$\{\ket{\psi_{j}}\}_{j=0}^{k}$ the non zero entries of the matrix $H$ are
$\bra{\psi_{j}}H\ket{\psi_{j\pm 1}}=1$ (17)
We now follow the idea in [Chi04] and modify the Hamiltonian in (15) in the following way:
$H^{\prime}:=\sum_{j=1}^{k}\sqrt{j(k+1-j)}\;H_{j}$ (18)
The idea in [Chi04] is to associate each state in {|$\{\ket{\psi_{j}}\}_{j=0}^{k}$ to a quantum system of total angular momentum $\frac{k}{2}(\frac{k}{2}-1)$ with $z$ component $j-\frac{k}{2}$. The association is possible by the fact that a system with total angular momentum $\frac{k}{2}(\frac{k}{2}-1)$ will allow $k+1$ states with $z$ components of values $-\frac{k}{2},-\frac{k}{2}+1,...,\frac{k}{2}-1,\frac{k}{2}$. These are exactly the $k+1$ states in {|$\{\ket{\psi_{j}}\}_{j=0}^{k}$ and it is possible to move among them defining the corresponding ladder operators :
$L_{-}\ket{\psi_{j}}$ $=\sqrt{j(k+1-j)}\ket{\psi_{j-1}}$ (19)
$L_{+}\ket{\psi_{j}}$ $=\sqrt{(k-j)(j+1)}\ket{\psi_{j+1}}$ (20)
By the algebra of the angular momentum, the $x$ component of the total angular momentum $J_{x}$ can be expressed as
$J_{x}=\frac{1}{2}(L_{+}+L_{-})$ (21)
Comparing Eq.(17) with Eq. (19) it is clear that $H^{\prime}$ is exactly the $x$ component of the angular momentum operator defined over the states {|$\{\ket{\psi_{j}}\}_{j=0}^{k}$. As $J_{x}$ rotates between the states with $z$ component ±$\pm\frac{k}{2}$ in time $t=\pi$, this concludes the proof.
A.1 Classical hardness
Theorem 4 (Lemma 1 in the main text).
Under the assumption that there exists a distribution $\mathcal{D}^{\prime}$ such that $\mathrm{BQP}\not\subset\mathrm{HeurP}^{\mathcal{D}^{\prime}}/\mathrm{poly}$, then for any BQP-complete language there exists an associated input distribution $\mathcal{D}$ and a Hamiltonian $H_{\mathrm{hard}}$ such that no randomized polynomial-time classical algorithm $A_{c}$ is such that $h_{cl}(.)=A_{c}(T^{\bm{\alpha}},.)$ satisfies the learning condition of Def. 1 for the concept class $\mathcal{F}^{H_{\mathrm{hard}},O}_{\mathrm{evolved}}$ with $x$ sampled from $\mathcal{D}$.
Proof.
The main idea of the proof is to show that if there exists a classical algorithm which learns by Def. 1 the time evolution problem for any local Hamiltonian, then such algorithm could decide any BQP language on average implying $\mathrm{BQP}\subseteq\mathrm{HeurP}^{\mathcal{D}}/\mathrm{poly}$, for any $\mathcal{D}$. As in Lemma 3 in [GD23] it is proved that, under widely believed assumptions, for any BQP complete language $\mathcal{L}$ there always exists a distribution $\mathcal{D}$ such that $(\mathcal{L},\mathcal{D})\not\subset\mathrm{HeurP}/\mathrm{poly}$, such classical algorithm cannot exist. Be $\{U^{n}_{BQP}\}_{n}$ a family of quantum circuit which decides the BQP complete language $\mathcal{L}$, one circuit per size. Precisely, that means that for any $x\in\{0,1\}^{n}$ measuring the first qubit in the computational basis on the state $U^{n}_{BQP}\ket{\bm{x}}$ will output 1 (or 0) with probability greater that 2/3 if $\bm{x}\in\mathcal{L}$ ($\bm{x}\not\in\mathcal{L}$). It is obvious that the the quantum model $f^{O^{\prime}}=\text{Tr}[\rho(\bm{x})O^{\prime}]$ correctly decides every $\bm{x}\in\mathcal{L}$ if $\rho(\bm{x})=U^{n}_{BQP}\ket{\bm{x}}\bra{\bm{x}}(U^{n}_{BQP})^{\dagger}$ and $O^{\prime}=Z\otimes\underbrace{I\otimes...\otimes I}_{n-1}$. In fact measuring the first qubit of $U^{n}_{BQP}\ket{\bm{x}}$:
  1. 1.
    For all $\bm{x}\in\mathcal{L}$, $f^{O^{\prime}}(\bm{x})=$ Pr[the output of $U^{n}_{BQP}$ applied on the input |$\ket{\bm{x}}$ is $1$] - Pr[the output of $U^{n}_{BQP}$ applied to the input |$\ket{\bm{x}}$ is 00] $\geq 2/3-1/3=1/3$.
  2. 2.
    For all $\bm{x}\notin\mathcal{L}$, $f^{O^{\prime}}(\bm{x})=$ Pr[the output of $U^{n}_{BQP}$ applied on the input |$\ket{\bm{x}}$ is $1$] - Pr[the output of $U^{n}_{BQP}$ applied to the input |$\ket{\bm{x}}$ is 00] $\leq 1/3-2/3=-1/3$.
Therefore, as $f^{O^{\prime}}(\bm{x})>0$ if $\bm{x}\in\mathcal{L}$ and $f^{O^{\prime}}(\bm{x})<0$ if $\bm{x}\notin\mathcal{L}$ such quantum model could efficiently decide the language $\mathcal{L}$. By Lemma 3 for every $n$ there exist a 4-local Hamiltonian $H_{\mathrm{\mathrm{hard}}}$ such that evolving the state |$\ket{\bm{x}}$ for a time $t=\pi$ under it will produce the state $U^{n}_{\mathrm{BQP}}\ket{\bm{x}}$ (on the work register). Since such constructed $H_{\mathrm{hard}}$ is local, its time evolution can be efficiently implemented on a quantum computer. Working towards a contradiction, assume now that there exists an algorithm $A_{c}$ which learns $\mathcal{F}^{H_{\mathrm{hard}},O}_{\mathrm{evolved}}$ with respect to the distribution $\mathcal{D}$ by Def.1. Then it means that there exists a randomized classical algorithm which makes use of data such that $h_{cl}(.)=A_{c}(H_{\mathrm{hard}},\mathcal{T}^{\bm{\alpha}},.)$ satisfies $\forall\epsilon\geq 0$ and $\forall f^{\bm{\alpha}}\in\mathcal{F}^{H_{\mathrm{hard}},O}_{\mathrm{evolved}}$ the condition:
$\mathbb{E}_{\bm{x}\sim\mathcal{D}}\big{[}\;|f^{{\bm{\alpha}}}(\bm{x})-h_{cl}(\bm{x})|^{2}\;\big{]}\leq\epsilon$ (22)
with high probability over the choice of the training dataset and the internal randomization of $A_{c}$. We want now to show that the classical model $h_{cl}$ would be able to decide the language $\mathcal{L}$ on average with respect to $\mathcal{D}$. Firstly notice that as the algorithm $A_{c}$ succeeds with high probability, there exists a fixed dataset $\mathcal{T}^{\prime\bm{\alpha}}$ and a fixed bitrsting $r^{\prime}$ for its internal randomization such that $h^{\prime}_{cl}(.)=A_{c}(H_{\mathrm{hard}},\mathcal{T}^{\prime\bm{\alpha}},r^{\prime},.)$ satisfies Eq.(22). Throughout the reminder of the proof let us fix such dataset $\mathcal{T}^{\prime\bm{\alpha}}$ and random bitstring $r^{\prime}$. Also, since Def. 1 assumes Eq. (22) is satisfied for every concept ${\bm{\alpha}}$, such an $A_{c}$ would be able to evaluate on avarage the particular concept ${\bm{\alpha}}^{\prime}$ which corresponds to the observable $O({\bm{\alpha}}^{\prime})=O^{\prime}=Z\otimes I\otimes...\otimes I$. We now show that such classical algorithm would be able to decide the language $\mathcal{L}$ in $\mathrm{HeurP}/\text{poly}$ with respect to the distribution $\mathcal{D}$. Remember that a language $L$ and distribution $D$ is in $\mathrm{HeurP}/\text{poly}$ if there exist a classical algorithm $\mathcal{A}$ and a polynomial-size advice string $\omega$ such that
$\text{Pr}_{\bm{x}\sim\mathcal{D}}[\mathcal{A}(\omega,\bm{x})\neq L(\bm{x})]\leq\epsilon$ (23)
where $L(\bm{x})=1$ if $\bm{x}\in L$ and 0 otherwise. Notice that Eq.(23) is equivalent to
$\sum_{\bm{x}\in\mathcal{X}_{err}}p(\bm{x})\leq\epsilon$ (24)
where $\mathcal{X}_{err}$ is the set of input $\bm{x}$ such that $[\mathcal{A}(\omega,\bm{x})\neq L(\bm{x})]$, i.e. $\mathcal{X}_{err}=\{\bm{x}\in\{0,1\}^{n}\;\;|\;\;[\mathcal{A}(\omega,x)\neq L(\bm{x})\}$, and $p(\bm{x})$ is the probability associated to the input $\bm{x}$ with respect to the distribution $\mathcal{D}$.
[…]
$\epsilon^{\prime}$ $\geq\mathbb{E}_{\bm{x}\sim\mathcal{D}}\big{[}\;|f^{{\bm{\alpha}}}(\bm{x})-h^{\prime}_{cl}(\bm{x})|^{2}\;\big{]}$ (25)
$\geq\sum_{\bm{x}}|f^{{\bm{\alpha}}}(\bm{x})-h^{\prime}_{cl}(\bm{x})|^{2}p(\bm{x})$ (26)
$\geq\sum_{\bm{x}\in\mathcal{X}}|f^{{\bm{\alpha}}}(\bm{x})-h^{\prime}_{cl}(\bm{x})|^{2}p(\bm{x})+\sum_{\bm{x}\not\in\mathcal{X}}|f^{{\bm{\alpha}}}(\bm{x})-h^{\prime}_{cl}(\bm{x})|^{2}p(\bm{x})$ (27)
$\geq 0.09\sum_{x\in\mathcal{X}}p(\bm{x})$ (28)
From last equation we obtain that $\sum_{\bm{x}\in\mathcal{X}}p(\bm{x})\leq\epsilon$. Notice that the classical algorithm $h^{\prime}_{cl}$ is such that $h^{\prime}_{cl}(\bm{x})>0$ if $x\in\mathcal{L}$ and $h^{\prime}_{cl}(\bm{x})<0$ if $\bm{x}\in\mathcal{L}$ for every $\bm{x}\in\{0,1\}^{n}/\mathcal{X}$. Then, the condition $\sum_{\bm{x}\in\mathcal{X}}p(\bm{x})\leq\epsilon$ is exactly the defining condition of $\mathrm{HeurP}/\mathrm{poly}$ in Eq.(24). This implies that the classical algorithm $A_{cl}$ using as advice $\mathcal{T}^{\prime\bm{\alpha}}$ and $r^{\prime}$ is able to decide the language $\mathcal{L}$ in $\mathrm{HeurP}/\mathrm{poly}$ with respect to the distribution $\mathcal{D}$. This concludes the proof. ∎
A.2 Quantum Learnability
In this Section we prove that Algorithm 1 satisfies the learning condition in Def. 1 using only polynomial resources in both time and samples. As a first step, notice that the quanutm states $\rho_{H_{\mathrm{hard}}}(\bm{x})$ in Eq. (1) of $\mathcal{F}^{H_{\mathrm{hard}},O}_{\mathrm{evolved}}$ are easily preparable on a quantum computer.
Since the quantum algorithm can efficiently prepare time evolved states of local Hamiltonians, the quantum states $\rho_{H_{\mathrm{hard}}}(\bm{x})=U\ket{\bm{x}}\bra{\bm{x}}U^{\dagger}$ with $U=e^{iH_{\mathrm{\mathrm{hard}}}\pi}$ can efficiently be prepared on a quantum computer. In order to demonstrate the quantum learnability of $\mathcal{F}^{H_{\mathrm{hard}},O}_{\mathrm{evolved}}$, all that remains is to rigorously bound the sample and time complexity of Algorithm 1 in the following Theorem 6. As the bound obtained on the sample complexity derives from the generalization bound of the LASSO algorithm [MRT18, LHT+24], we first repeat this result.
Theorem 5 (Theorem 11.16 in [MRT18]).
Let $\mathcal{X}\subseteq\mathbb{R}^{A}$ and $\mathcal{C}=\{\bm{x}\in\mathcal{X}\mapsto\vec{w}\cdot\bm{x}:\|\vec{w}\|_{1}\leq B\}$. Let $\mathcal{S}=((\bm{x}_{1},y_{1}),\dots,(\bm{x}_{N},y_{N}))\in(\mathcal{X}\times\mathcal{Y})^{N}$. Let $\mathcal{D}$ denote a distribution over $\mathcal{X}\times\mathcal{Y}$ according to which the training data $\mathcal{S}$ is drawn. Assume that there exists $r_{\infty}>0$ such that for all $\bm{x}\in\mathcal{X}$, $\|x\|_{\infty}\leq r_{\infty}$ and $M>0$ such that $|h(\bm{x})-y|\leq M$ for all $(\bm{x},y)\in\mathcal{X}\times\mathcal{Y}$. Then, for any $\delta>0$, with probability at least $1-\delta$, each of the following inequalities holds for all $h\in\mathcal{C}$:
$\mathbb{E}_{(\bm{x},y)\sim\mathcal{D}}[|h(\bm{x})-y|^{2}]:=R(h)\leq\hat{R}_{\mathcal{S}}(h)+2r_{\infty}BM\sqrt{2\log(2|A|)N}+M\sqrt{2\log(1/\delta)}/{2N}$
where $R(h)$ is the prediction error for the hypothesis $h$ and $\hat{R}_{\mathcal{S}}(h)$ is the training error of $h$ on the training set $\mathcal{S}$.
We are now ready to state Theorem 6 which provides a rigorous guarantee on the number of samples and the time complexity required by Algorithm 1.
Theorem 6 (Lemma 2 in the main text).
Given $n$, $\delta,\epsilon\geq 0$ and a training data set $\mathcal{T}^{\bm{\alpha}}_{N}=\{(\bm{x}_{\ell},y_{\ell})\}_{\ell=1}^{N}$ of size
$N=\mathcal{O}\left(\frac{\log(n/\delta)n}{\epsilon}\right)$ (29)
where $x_{\ell}$ is sampled from an unknown distribution $\mathcal{D}$ over $x\in\{0,1\}^{n}$ and $|y_{\ell}-\text{Tr}(O({\bm{\alpha}})\rho(\bm{x}_{\ell}))|\leq\epsilon$ for any geometrically local observable $O({\bm{\alpha}})=\sum_{i}^{m}{\bm{\alpha}}_{i}P_{i}$ with $P_{i}\in\{I,X,Y,Z\}^{\otimes n}$ and ${\bm{\alpha}}\in[-1,1]^{m}$, and $\rho(\bm{x})=U(\bm{x})\ket{0}\bra{0}U^{\dagger}(\bm{x})$. Then there exists a quantum algorithm $\mathcal{A}_{q}(\mathcal{T}_{N},x)=h(\bm{x})$ such that:
$\mathbb{E}_{x\sim\mathcal{D}}|h(\bm{x}_{l})-\text{Tr}[O({\bm{\alpha}})\rho(\bm{x})]|^{2}\leq\epsilon$ (30)
with probability at least $1-\delta$. The computational time of the quantum algorithm is bounded in $\mathcal{O}(nN)$
Proof.
The proof of the theorem is based on the well known bound on the prediction error of the LASSO algorithm of Theorem 5. Consider the algorithm described in Algorithm 1. First, we demonstrate now that Algorithm 1 satisfies the condition of Theorem 6. In our setting, the input space of the learned model $h$ is $\mathcal{X}=[-1,1]^{m}\subset\mathbb{R}^{m}$ as we consider $h$ a function of the $m-$dimensional feature vector $\phi(\bm{x})$. Clearly, $||\phi(\bm{x})||_{\infty}\leq 1$ for all $x\in\mathcal{X}$. The hypothesis class $\mathcal{C}$ of our algorithm is given by the set of function of the same form of the learned $h$, i.e. $\mathcal{C}=\{\phi(\bm{x})\in\mathcal{X}\rightarrow w\cdot\phi(\bm{x}):||w||_{1}\leq B\}$ with $B=poly(n)$. With respect to Theorem 6, we can also choose $M=poly(n)$ so that $|h(\bm{x}_{l})-y_{l}|<M$ for all $l=1,...,N$:
$|h(\bm{x}_{l})-y_{l}|\leq|w\cdot\phi(\bm{x}_{l})|+|y_{l}|\leq||w||_{1}||\phi(\bm{x}_{l})||_{\infty}+2\leq poly(n)$ (31)
Where the second inequalitiy follows by Hölder’s inequality. By Theorem 6 then, the bound on the prediction error $R(h)$ of the learned model $h(\bm{x})=w^{*}\phi(\bm{x})$ is
$R(h)\leq\hat{R}(h)+2BM\sqrt{\frac{2\log(2m)}{N}}+M^{2}\sqrt{\frac{\log\delta^{-1}}{2N}}$ (32)
where $\hat{R}(h)$ is the training error on the dataset $T^{\bm{\alpha}}_{N}$.
We now bound the training error $\hat{R}(h)$. Let $\epsilon_{1}$ be the maximum sampling error associated to the measurement of a Pauli observable on $\rho(\bm{x}_{\ell})$, i.e.
$\epsilon_{1}=\max_{\begin{subarray}{c}i\in[1,...,m]\\ \bm{x}\in\{0,1\}^{n}\end{subarray}}\;\;|[\rho(\bm{x})P_{i}]-\phi(\bm{x}_{i})|$ (33)
Let also $\epsilon_{2}$ the maximum sampling error associated to $y_{\ell}$, such that:
$|y_{\ell}-\text{Tr}[O({\bm{\alpha}})\rho(\bm{x}_{l})]|\leq\epsilon_{2}\;\;\;\;\;\forall(\bm{x}_{\ell},y_{\ell})\in T^{\bm{\alpha}}_{N}$ (34)
We can now derive a bound on the training error for the optimal value of $w_{opt}={\bm{\alpha}}$ of the model $h_{opt}(\bm{x})={\bm{\alpha}}\cdot\phi(\bm{x})$:
$\hat{R}(h_{opt})=\frac{1}{N}\sum_{\ell=1}^{N}|h_{opt}(\bm{x}_{\ell})-y_{\ell}|^{2}$ $\leq\max_{\ell}|h(\bm{x}_{\ell})-y_{\ell}|$ (35)
$\leq(|\text{Tr}[\rho(\bm{x}_{\ell^{*}})O({\bm{\alpha}})]-h(\bm{x}_{\ell^{*}})|+|\text{Tr}[\rho(\bm{x}_{\ell^{*}})O({\bm{\alpha}})]-y_{\ell}|)^{2}$ (36)
$\leq\left(\left(\sum_{i}|{\bm{\alpha}}_{i}|\right)\epsilon_{1}+\epsilon_{2}\right)^{2}$ (37)
$\leq(B\epsilon_{1}+\epsilon_{2})^{2}$ (38)
In practice, we can require to the LASSO algorithm to obtain a $w^{*}$ which achieves a training error at most $\epsilon_{3}/2$ larger than the optimal one. We can obtain such precision by setting $B=poly(n)$. Formally, we have
$\hat{R}(h)\leq\frac{\epsilon_{3}}{2}+\min_{\begin{subarray}{c}w\in\mathbb{R}^{m}\\ ||w||_{1}\leq B\end{subarray}}\frac{1}{N}\sum_{\ell=1}^{N}|w\cdot\phi(\bm{x}_{\ell})-\text{Tr}[\rho(\bm{x}_{\ell})O({\bm{\alpha}})]|^{2}$ (39)
Because we have set $B=poly(n)$, we have that the second term must be at most $\hat{R}(h_{opt})$ and therefore we have
$\hat{R}(h)\leq(B\epsilon_{1}+\epsilon_{2})^{2}+\frac{\epsilon_{3}}{2}$ (40)
We note that as $B\leq\mathcal{O}(poly(n))$ we can bound the error $\epsilon^{\prime}_{1}=B\epsilon_{1}$ to scale polynomially to zero by just reducing the sampling error $\epsilon_{1}$ using polynomially many more copies of each $\rho(\bm{x}_{l})$. Thus we can rewrite equation (32) as
$R(h)\leq(\epsilon^{\prime}_{1}+\epsilon_{2})^{2}+\frac{\epsilon_{3}}{2}+2BM\sqrt{\frac{2\log(2m)}{N}}+M^{2}\sqrt{\frac{\log(\delta)^{-1}}{2N}}$ (41)
Then, in order to bound the prediction error above by $\epsilon=(\epsilon^{\prime}_{1}+\epsilon_{2})^{2}+\epsilon_{3}$ we need to choose an N such that
$2BM\sqrt{\frac{2\log(2m)}{N}}+M^{2}\sqrt{\frac{\log(\delta)^{-1}}{2N}}\leq\frac{\epsilon_{3}}{2}$ (42)
By substituting $M=B$ we obtain that:
$2BM\sqrt{\frac{2\log(2m)}{N}}+M^{2}\sqrt{\frac{\log(\delta)^{-1}}{2N}}\leq\left(\frac{B^{2}}{\sqrt{N}}(\sqrt{2\log(2\mathcal{O}(n))}+\sqrt{\log(\delta^{-1})}\right)$ (43)
so that it is upper bounded by $\epsilon_{3}/2$ choosing $N$ as
$N$ $=2\frac{B^{4}\sqrt{2\log(\mathcal{O}(n)/\delta)}}{\epsilon_{3}}$ (44)
$\sim\frac{\log(n/\delta)poly(n)}{\epsilon_{3}}$ (45)
By setting $\epsilon^{\prime}_{1}=0.2\epsilon$, $\epsilon_{2}=\epsilon$, and $\epsilon_{3}=0.4$ we have $(\epsilon^{\prime}_{1}+\epsilon_{2})^{2}+\epsilon_{3}\leq\epsilon$ and recover the claim of the Theorem.
Finally we bound the efficiency of our algorithm. The training time is dominated by the creation of the feature map $\phi(\bm{x}_{\ell})$ for each training point and by the LASSO regression over the corresponding feature space. To create the vector $\phi(\bm{x})$ the quantum algorithm needs to prepare multiple copies of $\rho(\bm{x}_{\ell})$ $\forall\bm{x}_{\ell}\in T_{N}^{\bm{\alpha}}$. As seen before, only a polynomial number of copies are sufficient to achieve a desired error $\epsilon^{\prime}_{1}$, so that the whole process takes time $\mathcal{O}(nN)$. For the LASSO regression, it is known that to obtain a training error at most $\epsilon_{3}/2$ larger than the optimal function value, the LASSO algorithm on the feature space of $\phi(\bm{x})$ can be executed in time $\mathcal{O}\left(\frac{m_{\phi}\log m_{\phi}}{\epsilon^{2}_{3}}\right)$ [HK12], where in our case $m_{\phi}=m$. It is easy to show that even this time is bounded by $\mathcal{O}(nN)$. The prediction time corresponds to the time for the evaluation of the learned model $h(\bm{x})=w^{*}\cdot\phi(\bm{x})$ which takes time $\mathcal{O}(m)\sim\mathcal{O}(n)$. In conclusion, the overall time of the quantum algorithm is bounded by $\mathcal{O}(nN)$
Appendix B Proof of Theorem 3 and Corollary 1
We rewrite here Theorem 3 form the main text, formalizing the result and providing a detailed proof.
Theorem 7 (Theorem 3 in the main text).
Given a discrete set of polynomially describle quantum states $S=\{\ket{\psi_{\ell}}\}_{\ell}$ and measurements operators $Q=\{Q_{m}\}_{m}$. Then for every efficient (non-adaptive) algorithm $\mathcal{A}_{W}$ which learns an arbitrary unitary $W$ from query access to it on states from $S$ and measured by operators in $Q$ there exists a distribution $\mathcal{D}$ over $\{0,1\}^{n}$, a family of unitaries $\{U(\bm{x})\}_{\bm{x}}$ and a measurement $O$ for which the concept class $\mathcal{M}_{U,W,O}$ exhibits a learning advantage.
Proof.
We can think at the algorithm $\mathcal{A}_{W}$ as an algorithm which receives in input pairs of {(|$\{(\ket{\psi_{\ell}},y^{m_{\ell}}_{\ell})_{\ell}\}$, where |$\ket{\psi_{\ell}}\in S$ and $y^{m_{\ell}}_{\ell}\in\mathbb{R}$ is the measurement outcome of a randomly selected operator $Q_{m_{l}}\in Q$ on |$\ket{\psi_{\ell}}$, and for any $\epsilon\geq 0$ outputs a matrix $\tilde{W}$ such that $||\tilde{W}-W||\leq\epsilon$ in some norm distance. The idea to construct a $\mathcal{M}_{U,W,O}$ which exhibits a separation is the following, illustrated also in Figure 1. Consider the family of quantum states {|$\{\ket{\psi_{U}(\bm{x})}\}_{\bm{x}}$ constructed by parameterized unitaries $\{U(\bm{x})\}_{\bm{x}}$ in the following way, depending on the input $\bm{x}\in\{0,1\}^{n}$:
  • The first $1+n_{Q}$ qubit of |$\ket{\psi_{U}(\bm{x})}$ are initializated in the state |$\ket{x_{1}}\otimes\ket{x_{2}x_{3}...x_{n_{Q}+1}}$, the remaining $n_{S}=n-(1+n_{Q})$ qubit $\bm{x_{S}}=x_{n-n_{S}}x_{n-(n_{S}-1)}...x_{n}$ are in the state |$\ket{0^{\otimes n_{S}}}$.
  • If the first bit $x_{1}$ is 0, then $U(\bm{x})$ prepares on the $n_{S}$ qubits register the state |$\ket{\psi_{\bm{x_{S}}}}$ from the set $S$ described by the bitstring $\bm{x_{S}}\in\{0,1\}^{n_{S}}$. As we assumed that every state in $S$ allows a polynomial classical description, we only need an input $\bm{x}\in\{0,1\}^{n}$ of polyinomial size. We thus have:
[…]
  • If the first bit $x_{1}$ is 1, then on the $n_{S}$ qubit register $U(\bm{x})$ prepares the state |$\ket{\psi_{BQP}(\bm{x_{S}})}$ such that $\bra{\psi_{BQP}(\bm{x_{S}})}Z\otimes I\otimes...\otimes I\ket{\psi_{BQP}(\bm{x_{S}})}$ outputs $+1/3$ if the $n_{S}$ bitstring $\bm{x_{S}}$ belongs to an arbitrary (previously fixed) BQP complete language $\mathcal{L}$ defined over input $\bm{x_{S}}\in\{0,1\}^{n_{S}}$, while it outputs $-1/3$ otherwise. Following the same arguments used in the proof of Lemma 1, for any $\mathcal{L}\in\mathrm{BQP}$ it always exists such $U(\bm{x})$. We thus have:
[…]
Regarding the unknown observable $O({\bm{\alpha}})$, we will consider a measurement operator of the kind Eq.(9). Specifically we define a controlled operator $V_{A}$, controlled by the $1+n_{Q}$ qubit register |$\ket{x_{1}x_{2}x_{3}...x_{n_{Q}+1}}$, such that $V_{A}=\sum_{0\bm{x_{Q}}}\ket{0\bm{x_{Q}}}\bra{0\bm{x_{Q}}}\otimes V(\bm{x_{Q}})$ with $\bm{x_{Q}}=x_{2}x_{3}...x_{n_{Q}+1}$ if $x_{1}=0$ and $V_{A}=\sum_{1\bm{x_{Q}}}\ket{1\bm{x_{Q}}}\bra{1\bm{x_{Q}}}\otimes I^{\otimes n_{S}}$ if $x_{1}=1$. The unitary matrices $V(\bm{x_{Q}})$ are defined such that they rotate the measurement $O$ into the observable $Q_{\bm{x_{Q}}}\in Q$ described by the bitstring $\bm{x_{Q}}=x_{2}x_{3}...x_{n_{Q}+1}$, i.e. $Q_{\bm{x_{Q}}}=V(\bm{x_{Q}})OV^{\dagger}(\bm{x_{Q}})$ where $O$ could be taken as the local observable $Z_{1}=Z\otimes I^{\otimes n_{S}-1}$ on the $n_{S}$ qubit register. The final observable $O(\bm{\alpha})$ will then be $O(\bm{\alpha})=(I^{\otimes(1+n_{Q})}\otimes W(\bm{\alpha})\;V_{A})O(V_{A}\;I^{\otimes(1+n_{Q})}\otimes W(\bm{\alpha}))^{\dagger}$, with $W({\bm{\alpha}})$ an arbitrary unitary parameterized by ${\bm{\alpha}}$. Furthermore, we define $W({\bm{\alpha}}=\vec{0})=I\otimes...\otimes I$. We can now construct a distribution $\mathcal{D}$ such that the concept class $\mathcal{M}_{U,W,O}$ exhibits a learning separation. For Lemma 1, there exists a distribution $\mathcal{D}_{\mathrm{hard}}$ for which the concept $f^{0}=\text{Tr}[\rho_{U}(\mathbf{x})V_{A}OV^{\dagger}_{A}]$ cannot be efficiently learned by a classical algorithm when $U(\bm{x})$ is the circuit which decides the BQP-complete language $\mathcal{L}$ over the $n_{S}$-sized bitstrings $\bm{x_{S}}$. We now define the following distribution $\mathcal{D}$ on the input bitsrings $\bm{x}\in\{0,1\}^{n}$:
  • The first bit $x_{1}$ of $\bm{x}$ is extracted randomly with equal probability between 0 or 1.
  • If $x_{1}=0$ then the other $n-1$ bit $x_{2}x_{3}...x_{n}$ are extracted from the required distribution by $\mathcal{A}_{W}$ to learn the unitary $W(\bm{\alpha})$.
  • If $x_{1}=1$ then the following $n_{Q}$ bit $\bm{x_{Q}}$ are extracted uniformly at random while the $n_{S}$ bit $\bm{x_{S}}$ are sampled with respect to the distribution $\mathcal{D}_{\mathrm{hard}}$.
Classical hardness The classical hardness of the learning task comes directly from the proof of Lemma 1. Consider the concept $f^{0}\in\mathcal{M}_{U,W,M}$ defined by $W(\bm{0})=I^{\otimes n_{S}}$. For the same reasoning of the proof of Lemma 1 there cannot exists a polynomial sized classical circuit which evaluates $f^{0}(\bm{x})$ correctly on average with respect to the distribution $\mathcal{D}$ for the cases where $x_{1}=1$. As $x_{1}$ is equally sampled between 0 and 1, no classical algorithm can meet the learning condition of Eq. 1 on half of the input bitstrings, thus it can not learn the concept $f^{0}\in\mathcal{M}_{U,W,O}$ in polynomial time for every $\epsilon$. As the learning algorithm must suceed for every $\bm{\alpha}$, this suffices to prove the classical hardness of the learning task.
Quantum learnability Recall that the data the learning algorithm receives for a concept $f^{\bm{\alpha}}\in\mathcal{M}_{U,W,M}$ are $\mathcal{T}^{\bm{\alpha}}=\{\bm{x}_{\ell},y_{\ell}\}_{\ell}$ with $\mathbb{E}[y_{\ell}]=\text{Tr}[\rho_{U}(\bm{x}_{\ell})O({\bm{\alpha}})]$ and $\rho_{U}(\bm{x}_{\ell})=\ket{\psi_{U}(\bm{x})}\bra{\psi_{U}(\bm{x})}$ |. Now, in the case the first bit of $\bm{x}_{\ell}$ is 0, $\mathcal{T}^{\bm{\alpha}}$ are exaclty the pairs {(|$\{(\ket{\psi_{\ell}},y^{m}_{\ell})\}_{\ell}$ required by the algorithm $\mathcal{A}_{W}$ to learn $W(\bm{\alpha})$ and thus $O(\bm{\alpha})$. As $x_{1}=0$ for half of the training samples in $\mathcal{T}^{\bm{\alpha}}$, the quantum algorithm is able to learn $O(\bm{\alpha})$ and evaluate it on any input state.
We provide here the the rigorous proof of Corollary 1. The proof follows the same steps of the one for Theorem 3 above while concretizing the result for shallows $W$.
[…]
Lemma 4 (Lemma 10 in [HLB+24], Learning a few-body observable with an unknown support).
Given an error $\epsilon$, failure probability $\delta$, an unknown $n$-qubit observable $O$ with $\|O\|_{\infty}\leq 1$ that acts on an unknown set of $k$ qubits, and a dataset $\mathcal{T}_{O}(N)=\{|\psi_{\ell}\rangle=\bigotimes_{i=1}^{n}|\psi_{\ell,i}\rangle,v_{\ell}\}_{\ell=1}^{N}$, where $|\psi_{\ell,i}\rangle$ is sampled uniformly from stab1 and $v_{\ell}$ is a random variable with $\text{E}[v_{\ell}]=\langle\psi_{\ell}|O|\psi_{\ell}\rangle$, $|v_{\ell}|=\mathcal{O}(1)$. Given a dataset size of
$N=\frac{2^{\mathcal{O}(k)}\log(n/\delta)}{\epsilon^{2}},$ (48)
with probability at least $1-\delta$, we can learn an observable $O^{\prime}$ such that $\|O^{\prime}-O\|_{\infty}\leq\epsilon$ and $\textup{supp}(O^{\prime})\subseteq\textup{supp}(O)$. The computational complexity is $\mathcal{O}(n^{k}\log(n/\delta)/\epsilon^{2})$.
The proof of Lemma 1 then goes by explicitly constructing a family of parametrized circuit $\{U(\bm{x})\}_{\bm{x}}$ and an input distribution $\mathcal{D}$ such that the concept class $\mathcal{M}_{U,W,M}$ is quantum learnable using Lemma 4 while still being classically hard. Let rewrite here Corollary 1
Corollary 2 (Corollary 1 in the main text).
There exists a family of parametrized unitaries $\{U(\bm{x})\}_{\bm{x}}$ and parametrized shallow circuits $\{W({\bm{\alpha}})\}_{\bm{\alpha}}$, a measurement $O$ and a distribution $\mathcal{D}$ over $\bm{x}\in\{0,1\}^{n}$ such that the concept class $\mathcal{M}_{U,W,O}$ is not classically learnable with respect to the input distribution $\mathcal{D}$. However, there exists a quantum algorithm which learns $\mathcal{M}_{U,W,O}$ on the input distribution $\mathcal{D}$.
Proof.
Let first define the set of unitaries $\{U(\bm{x})\}_{\bm{x}}$, $\{W(\bm{\alpha})\}_{\bm{\alpha}}$, measurement operator $O$ and input distribution $\mathcal{D}$ on which the separation result holds. We provided a graphical representation of it in Figure 2. Be $n_{S}=\left\lfloor{\frac{n}{3}}\right\rfloor$, we define the set $\{U(\bm{x})\}_{\bm{x}}$ as the set of unitaries which act on a $n_{S}+1$ qubit system in the following way, depending on the first bit $x_{1}$ of the input $\bm{x}$.
If $x_{1}=0$ then each $U(\bm{x})$ is defined such that
$U(\bm{x})\left(\ket{x_{1}}\otimes\ket{0^{n_{S}}}\right)=\ket{x_{1}}\otimes\ket{\psi_{stab}(\bm{x})}$ (49)
[…]
Let consider now the case when $x_{1}=1$. Be $\mathcal{L}$ a BQP-complete language defined over the input $\tilde{\bm{x}}\in\{0,1\}^{n_{S}}$. Then we define the set of $\{U(\bm{x})\}_{\bm{x}}$ in the following way:
$U(\bm{x})\left(\ket{x_{1}}\otimes\ket{0^{n_{S}}}\right)=\ket{x_{1}}\otimes\ket{\psi_{BQP}(\bm{x})}$ (50)
where |$\ket{\psi_{BQP}(\bm{x})}$ is such that $\bra{\psi_{BQP}(\bm{x})}Z\otimes I\otimes...\otimes I\ket{\psi_{BQP}(\bm{x})}=1/3$ if $\tilde{\bm{x}}\in\mathcal{L}$ and $-1/3$ if $\tilde{\bm{x}}\not\in\mathcal{L}$. As we described in the Proof of Lemma 1, such a circuit $U_{BQP}(\bm{x})$ always exists for any BQP language.
As a set of unitaries $\{W(\bm{\alpha})\}_{\bm{\alpha}}$ we consider any set of parametrized unitaries acting on the $n_{S}$ qubit register such that for each $\bm{\alpha}$ parameter $W(\bm{\alpha})$ is shallow and such that $W(\bm{0})=I^{\otimes n_{S}}$. The $n_{S}$ qubit register is then measured by the observable $M=Z\otimes I\otimes...\otimes I$. Finally we define the distribution $\mathcal{D}$ on the input bitstrings $\bm{x}\in\{0,1\}^{n}$ to be the following:
  • The first bit $x_{1}\in\{0,1\}$ of $\bm{x}$ is extracted randomly with equal probability.
  • If $x_{1}=0$ then the other $n-1$ bit $x_{2}x_{3}...x_{n}$ are extracted following the uniform distribution over $\{0,1\}^{n-1}$.
  • If $x_{1}=1$ then the following $n_{S}$ bits $x_{2}x_{3}...x_{n_{S}+1}$ are extracted following the distribution $\mathcal{D}_{\mathrm{hard}}$ such that $(\mathcal{L},\mathcal{D}_{\mathrm{hard}})\not\subset\mathrm{HeurP}/\mathrm{poly}$. As we assumed $\mathcal{L}\in\mathrm{BQP}$-complete such a distribution always exists by Lemma 3 in[GD23]. The other $n-(n_{S}+1)$ input bits are sampled uniformly at random.
We now show that the concept class $\mathcal{M}_{U,W,M}$ defined in Eq. 9 with $\{U(\bm{x})\}_{\bm{x}}$, $\{W(\bm{\alpha})\}_{\bm{\alpha}}$ and $M$ considered above exhibits a learning separation with respect the input distribution $\mathcal{D}$.
Classical hardness The argument is exactly the same as the one presented before in the proof of Theorem 3.
Quantum learnability The quantum learnability is guaranteed by Lemma 4 in [HLB+24]. Recall that the training data the learning algorithm receives for a concept $f^{\bm{\alpha}}\in\mathcal{M}_{U,W,M}$ are $\mathcal{T}^{\bm{\alpha}}=\{\bm{x}_{\ell},y_{\ell}\}_{\ell}$ with $\mathbb{E}[y_{\ell}]=\text{Tr}[\rho_{U}(\bm{x}_{\ell})O({\bm{\alpha}})]$. Now, in the case the first bit of $\bm{x}_{\ell}$ is 0, $\mathcal{T}^{\bm{\alpha}}$ is exactly the training set $\mathcal{T}_{\bm{\alpha}}$ required by the algorithm in Lemma 4 to learn $O(\bm{\alpha})$. Since for every $\bm{\alpha}$ the unitary $W(\bm{\alpha})$ is of shallow depth, the locality of $O(\bm{\alpha})$ scales logarithmic with the number of qubit $n_{S}$. Then Lemma 4 guarantees that the learning algorithm runs in polynomial time requiring a polynomial-sized dataset, a condition met as half of the training samples in $\mathcal{T}^{\bm{\alpha}}$ will suffice.
{quantikz}
[…]
Figure 2: Quantum model which exhibits a learning separation for the concept class $\mathcal{M}_{U,W,O}$ when $W$ is a shallow circuit. The unitary $U(\bm{x})$ prepares the state |$\ket{\psi_{U}(\bm{x})}=\ket{x_{1}}\otimes\ket{\psi^{n_{S}}_{U}(\bm{x_{S}})}$ where the form of |$\ket{\psi^{n_{S}}_{U}(\bm{x_{S}})}$ depends on $x_{1}$, the first bit of each input $\bm{x}\in\{0,1\}^{n}$. If $x_{1}=0$, then |$\ket{\psi^{n_{S}}_{U}(\bm{x_{S}})}=\ket{\psi_{stab}(\bm{x_{S}})}=\bigotimes_{i=1}^{n_{S}}\ket{\psi^{i}_{stab1}(\bm{x_{S}})}$ is the $n_{S}=\left\lfloor{\frac{n}{3}}\right\rfloor$ qubit tensor product of single-qubit stabilizers described by the classical bitstring $x_{2}x_{3}...x_{n}$. If $x_{1}=1$, then |$\ket{\psi^{n_{S}}_{U}(\bm{x_{S}})}$ is the quantum state which decides the $n_{S}$ input bits $\bm{x_{S}}=x_{2}x_{3}...x_{n_{S}+1}$, considered as input of a BQP-complete language $\mathcal{L}$ over the bitstrings $\bm{x_{S}}\in\{0,1\}^{n_{S}}$ $W({\bm{\alpha}})$ is a parametrized shallow unitary, to prove classical hardness it is sufficient to consider $W({\bm{\alpha}}=0)=I^{\otimes n_{S}}$ and the measurement operator to be $O=Z\otimes I\otimes...\otimes I$ on the $n_{S}$ qubits register.